OpenStudy (anonymous):

what does f '(1) stand for ? i know f(1) just replaces the x in a equation... but what does the apostrophe in f '(1) mean?.....

OpenStudy (anonymous):

it is inverse

OpenStudy (anonymous):

replace x=1 in the inverse function of f

OpenStudy (anonymous):

hmmm?? can u explain to me like in this equation?? x^2 - x +2

OpenStudy (anonymous):

it is a bit difficult to explain through this but i will try

OpenStudy (anonymous):

k ty =/

OpenStudy (anonymous):

i will type the whole question... in case... let f(x) = x^2 + ax + b. if f(1) = 6 and if f'(1)=-4, find the values of a and b and hence, find the local minimum point on y = f(x).

OpenStudy (amistre64):

f'(1) ususally denotes the derivative when x=1

OpenStudy (anonymous):

so f'(1) is just the derivitive of x^2 + ax + b?

OpenStudy (amistre64):

f' is teh derivative of that, the (1) means you replace x by 1 as usual

OpenStudy (anonymous):

ah ok, thx a lot ^.^

OpenStudy (amistre64):

yw